During an automobile sale, cars are selling at the rate of 12/x + 1 cars per day, where x is the number of days since the sale began. How many cars will be sold during the first 7 days of the sale?
1
Expert's answer
2014-03-31T11:06:34-0400
To find the number of sold cars during 7 days we should find the following sum
(12/1+1)+(12/2+1)+(12/3+1)+(12/4+1)+(12/5+1)+(12/6+1)+(12/7+1)=13+7+5+4+1+12/5+3+12/7+1=34+(84+60)/35 approx.=38